A motorist is speeding along at 120 km/hr when he sees a squirrel on the road 200 meters in front of him. He tries to stop, but it takes 12 seconds for his car to stop. Assume acceleration was constant. (i) What is the acceleration of the car?


1
Expert's answer
2021-05-17T09:59:39-0400

"s=vt-\frac{at^2}{2},"

"a=\frac{2vt-2s}{t^2},"

"a=2.8~\frac{m}{s^2}."